The density of unknown gas A is 2.0 g/L at 2oC and 1 atm. What its density in g/mL at 90 kPa and 24oC, given that 1 atm = 101.3 kPa?
Answer #1

Let, the molecular weight of the gas is M g/mole.

Given that the density of the gas is 2.0 g/L

Now, according to ideal gas equation density of the gas

p = PM/RT ( p= density , P = pressure, T= temperature)

Or, p1/p2 = P1 T2/P2 T1 (p1= initial density 2.0g/L =0.002g/mL, P1= initial pressure, 1atm= 101.3 kPa , T1=initial temperature 2°C = (273+2 ) =275K,P2 = final pressure = 90kPa, T2= final temperature =(273+24)=297K

Or, p2 = p1*P2*T1 / P1T2

   = 0.002* 90*275/ 101.3*297 g/mL

= 0.001645g/mL